Fermi National Laboratory
 

Operation Summary for April 23, 2010 - April 26, 2010

 

The Integrated Luminosity for the period from 4/19/10 to 4/26/10 was 51.99 pb-1.
NuMI reported receiving 8.69E18 protons on target during this same period.

Friday April 23

The day shift began with Operations monitoring the Tevatron (TeV) store 7758 (131.38E30) and the Recycler stash (230.4E10), and the Antiproton Source (Pbar) stacking (21.7E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.

During the day shift, at 1 PM, Operators prepared the Pbar AP0 area for an access into the water cage so that Pbar personnel could switch dump LCW lines - a period of NuMI-only studies took place during this time. The access ended and stacking resumed at 3:30 PM.

During the evening shift, Operations terminated store 7758 at 6:49 PM, put the TeV into shot setup, and established store 7759 at 8:08 PM with an average initial luminosity (AIL) of 368.8E30. Water group personnel brought the CUB column #3 online and opened the isolation valves. Operations reported that the MI power supply interlock loop tripped when the MI-30 quadrupole VCB opened up; everything reset. The chilled water booster pump for the computer room AC system tripped off. FESS Duty Personnel brought the backup pump on and temperatures stabilized.

Saturday April 24

The midnight shift began with Operations monitoring store 7759 (178.62E30) and the stash (115.6E10), with Pbar stacking (24.3E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, everything ran smoothly.

During the day shift, Operations terminated store 7759 at 12:53 PM, put the TeV into shot setup, and established store 7760 at 2:26 PM with an AIL of 376.2E30. Operators tuned the Linac and Booster. Operations reported that ODH alarms posted for several areas. They investigated and found them to be due to the dropping barometric pressure.

During the evening shift, Operations reported at 9:16 PM that store 7760 aborted due to a TeV ground fault indication on quadrupole magnet (C:DQ0). It was a clean abort with no quench. Operators investigated and discovered an ice-ball on the conning tower that caused the ground fault; operators melted the ice. A Pbar Accumulator quadrupole power supply (A:LQ) tripped off; the stack was lost. Operators investigated and discovered that one of the phases on the internal 480 VAC breaker in the supply suffered a broken bolt and came disconnected. Operations called in a power supply person.

Sunday April 25

The midnight shift began with Operations monitoring the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

During the midnight shift, stacking resumed at 12:41 AM. Operations recovered the TeV and put it into shot setup. Operations established store 7764 at 7:39 AM with an AIL of 338.3E30.

During the day shift, Operations reported that only minor downtime took place during shift. Operators tuned the Linac and Booster.

During the evening shift, Operations terminated store 7764 at 9:36 PM and prepared CDF and D0 for accesses into their collision halls. Operations put the TeV through a dry squeeze.

Monday April 26

The midnight shift began with Operations monitoring the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations put the TeV into shot setup at 1:11 AM and established store 7768 at 2:14 AM with an AIL of 381.8E30.

The Plans for Monday

The plans for today are to stack, stash, and store.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 04/28/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory